OpenStudy (anonymous):

Perform the following division (find the quotient). (36x^2 − 9x^3 −27x^4) /(−27x3)

OpenStudy (anonymous):

oh hi dere

OpenStudy (anonymous):

Using LONG DIVISION METHOD we first rearrange the terms in increasing order of their power and then divide x + 1/3 _________________________________ −27x^3 | −27x^4 −9x^3 + 36x^2 -------- 27x^4 ------------------------- -9x^3 + 36x^2 9x^3 ----------------- 36x^2

OpenStudy (anonymous):

no its wrong sir

OpenStudy (anonymous):

what answer do u hv ??

OpenStudy (anonymous):

oh crap i mistyped one second

OpenStudy (anonymous):

if the answer is x +1/3 then the computer said it was wrong :(

OpenStudy (anonymous):

pls check if u hv typed the question right or is there a mistake....

OpenStudy (anonymous):

yea i wrote it correctly

OpenStudy (anonymous):

its correct

OpenStudy (anonymous):

dont know what to tell ya guys. thanks for trying though

OpenStudy (anonymous):

36x^2 − 9x^3 −27x^4 -36x^2 + 9x^3 + 27x^4 -------------------- = ------------------------ −27x^3 27x^3 -36x^2 + 9x^3 + 27x^4 -36x^2 9x^3 27x^4 ------------------------ = -------- + -------- + -------- 27x^3 27x^3 27x^3 27x^3 -4 1 = ----- + ----- + x 3x 3 try this answer...

OpenStudy (anonymous):

yea thats it. thanks brosef

OpenStudy (anonymous):

welcome, as they say there's more than one way of skinning a cat...☺

OpenStudy (anonymous):

Oi! i have a cat :P

OpenStudy (anonymous):

don't worry, a cat has nine lives!!!!!☺